OpenStudy (anonymous):

find dy/dx y=2/x when x=2

OpenStudy (lgbasallote):

do you mean \[\frac{dy}{dx} + y = \frac 2x\] ??

OpenStudy (anonymous):

y=2/x a) find dy/dx when x=2

jimthompson5910 (jim_thompson5910):

y = 2/x y = 2x^(-1) Now use the power rule

OpenStudy (lgbasallote):

oh lol

OpenStudy (lgbasallote):

i thought it was a differential equation

OpenStudy (anonymous):

its been a while, sorry but whats the power rule?

jimthompson5910 (jim_thompson5910):

y = x^n dy/dx = nx^(n-1)